OpenStudy (anonymous):

Y=3x+2 Y=x+12 Using the substitution method can someone help me?

OpenStudy (texaschic101):

you know that y = 3x + 2.....so just sub 3x + 2 in for y in the other equation.. y = x + 12 3x + 2 = x + 12 (subtract x from both sides) 3x - x + 2 = x - x + 12 (combine like terms) 2x + 2 = 12 (now subtract 2 from each side) 2x + 2 - 2 = 12 - 2 (combine like terms) 2x = 10 (now divide by 2 to get x by itself) 2x/2 = 10/2 x = 5 now sub 5 in for x in either of the original equations y = x + 12 y = 5 + 12 y = 17 check... y = 3x + 2 17 = 3(5) + 2 17 = 15 + 2 17 = 17 (correct) ANSWER : x = 5 and y = 17 or (5,17) Do you have any questions ? If so, please ask.

OpenStudy (anonymous):

First solve the top equation for x, and put the result in the x of the 2nd equation. Looks like texa's gotcha covered!

OpenStudy (anonymous):

Thanks for helping. I'll have another question up in a little bit. @texaschic101

OpenStudy (texaschic101):

I probably shouldn't have given the answer...but everyone needs a break sometimes :)

OpenStudy (anonymous):

it really helped though i understand the steps and what to do. @texaschic101

OpenStudy (texaschic101):

thats good....but remember, if you have any question or concerns, please ask.

OpenStudy (anonymous):

3x+y=10 y=x-2 you think you could help me out @texaschic101

OpenStudy (texaschic101):

you know that y = x - 2. So sub x - 2 in for y in the first equation. Can you do that ?

OpenStudy (texaschic101):

I will walk you through this one.....I will not give the answer....you will find the answer. Let me start you off... 3x + y = 10 3x + x - 2 = 10 now get the x's on one side and everything else on the other side. Now it is your turn.

OpenStudy (texaschic101):

are you having trouble ?

OpenStudy (texaschic101):

3x + x - 2 = 10 (combine like terms then solve for x) let me know if you need help.....